OpenStudy (anonymous):

(Three Variable Systems with Elimination) Solve each system by elimination. Check your answers. 6q - r + 2s = 8 2q + 3r - s = -9 4q + 2r + 5s = 1 I just learned how to do this today in class. I pretty much understand how to do it, but I think I made a silly mistake somewhere. I started getting s = -21/17 and I was like "I don't think this is right..." Thanks SOOOO much in advance!! :)

OpenStudy (anonymous):

q= 1/2 r = -3 s =1

OpenStudy (anonymous):

@surdawi You need to explain it. That doesn't help

OpenStudy (anonymous):

@ChmE do you want to explain it or you want me to

OpenStudy (anonymous):

You go ahead

OpenStudy (anonymous):

6q - r + 2s = 8 ===> r = 6q +2s -8 2q + 3r - s = -9 ====> s =2q +3r +9 ===> s =2q +3(6q +2s -8) +9 ===> s=-4q+3 4q + 2r + 5s = 1 4q + 2(6q +2(-4q+3) -8) + 5(-4q+3) = 1 11 - 20q = 1 q=1/2 now plug in s=-4q+3 s=-4*1/2+3=1 r = 6q +2s -8 r=6*1/2+2*1-8 r=-3

OpenStudy (anonymous):

That is the substitution method. The question is asking for elimination.

OpenStudy (anonymous):

damn i did it using matrices at the beginning

OpenStudy (anonymous):

Yeah I can only use elimination haha! :)

OpenStudy (anonymous):

@ChmE your turn bro

OpenStudy (anonymous):

1) 6q - r + 2s = 8 2) 2q + 3r - s = -9 3) 4q + 2r + 5s = 1 STEP 1 Lets use 1) and 2) first 1) 6q - r + 2s = 8 2) (2q + 3r - s = -9)2 6q - r + 2s = 8 4q + 6r - 2s = -18 Add them 10q + 5r = -10

OpenStudy (anonymous):

STEP 2 Eliminate the same variable 1) (6q - r + 2s = 8)5 3) (4q + 2r + 5s = 1)-2 1) 30q - 5r +10s = 40 3) -8q + -4r + -10s = -2 Add them 22q - 9r = 38

OpenStudy (anonymous):

STEP 3 Put our new equations with only 2 variables together in one system A) (10q + 5r = -10)9 B) (22q - 9r = 38)5 A) 90q + 45r = -90 B)110q - 45r = 190 add them 200q = 100 q = 1/2

OpenStudy (anonymous):

STEP 4 Now that we know q we can plug it in to equation A) or B) to solve for r. STEP 5 Now that we have q and r, we can plug them into 1), 2), or 3) to get s

OpenStudy (anonymous):

Did that make sense?

OpenStudy (anonymous):

For Three Variable Systems with Elimination eliminate a variable in 1 & 2 and the same in 2 & 3. It should work for 3 & 1 and 3 & 2 as well. You just have to do double elimination, then eliminate one last time with your 2 new equations. It's a pain

OpenStudy (anonymous):

@ChmE thank you SOOOOOOO much!!!! Really helped a ton :)

OpenStudy (anonymous):

It really is a pain, it takes FOREVER!

OpenStudy (anonymous):

@ChmE Wait I can't do what you said, I have to make the first two equations cancel out a variable. For example, 1) (6q - r + 2s = 8) 3) (4q + 2r + 5s = 1) I would have to multiply 2 by like 2 to make it 12q - 2r + 4s = 16 so it would 2r would cancel out when I added +2r to it from the other equation.